In an auditorium the first row has 20 seats, the second row has 25 seats and so on. Write the algebric equation to find out how many seats will be in the twentieth row.
I figured out the answer of 115 but I have no idea how to write the equation.
Thank you
Riley

Answer by solver91311(24713)


There are two (immediately obvious) ways to do this.

Let n be the row number.

Then if row 1 has 20, row 2 has 20 + 5(1), and row 3 has 20 + 5 + 5 = 20 + 5(2), you can say that row n has 20 plus 5 times (the row number minus 1).

So:



Let's check for row 20:



The other way to do it is to consider an imaginary row 0 that has 5 less seats than the first row, namely 15. Then the first row has 15 plus 5(1) = 20, the second row has 15 + 5(2) = 25, and the nth row has 15 + 5n



Again, let's try row 20:



Same correct result.



John